# VA-GS: Enhancing the Geometric Representation of Gaussian Splatting via View Alignment

### v1

@@ Appendix C More Ablation Results
(2) Gaussian Initialization. Similar to prior works, we use COLMAP [36] to generate a sparse point cloud from the input images, which serves as the initialization for our 3D Gaussians. To evaluate the impact of initialization quality, we experiment with randomly generated point clouds using two alternative strategies: (i) uniform sampling within a cube, and (ii) sampling points on a spherical surface. As shown in Table 7, both alternatives degrade the final reconstruction quality compared to COLMAP initialization. However, the spherical initialization performs better than the cube sampling, likely due to its more uniform coverage and approximate enclosure of the scene geometry.
@@ Appendix E Limitations and Future Works
Despite achieving strong performance in both surface reconstruction and novel view synthesis, our method has several limitations that suggest promising directions for future work. First, our model assumes known camera poses as input, which may not always be available in real-world scenarios, especially when the number of input views is extremely sparse. Removing this requirement by exploring pose-free approaches [50, 59] is a compelling future direction that would increase the applicability of our method in unconstrained settings. Moreover, our method assumes a moderate number of input views. For extremely sparse-view settings, the multi-view consistency constraints may become less effective, reducing reconstruction quality. Second, the number of Gaussians grows significantly with the number of input views, potentially limiting the scalability of our method in large-scale or densely captured scenes. Designing more compact or adaptive Gaussian representations, or incorporating efficient pruning and regularization strategies, could help improve scalability without compromising reconstruction quality. Third, our method currently relies on a pre-trained feature extraction network to guide geometric alignment and ensure robustness under varying lighting and viewpoints. This dependence introduces limitations, particularly when the pre-trained features are suboptimal for specific domains or scenes. Exploring more reliable, possibly self-supervised, feature learning strategies or reducing reliance on external networks altogether would further improve robustness and generalization. Addressing these limitations can pave the way for more flexible, efficient, and generalizable Gaussian-based surface reconstruction frameworks.

### v2

@@ Appendix C More Ablation Results
(2) Gaussian Initialization. Similar to prior works, we use COLMAP [schoenberger2016sfm] to generate a sparse point cloud from the input images, which serves as the initialization for our 3D Gaussians. To evaluate the impact of initialization quality, we experiment with randomly generated point clouds using two alternative strategies: (i) uniform sampling within a cube, and (ii) sampling points on a spherical surface. As shown in Table 7, both alternatives degrade the final reconstruction quality compared to COLMAP initialization. However, the spherical initialization performs better than the cube sampling, likely due to its more uniform coverage and approximate enclosure of the scene geometry.
@@ Appendix E Limitations and Future Works
Despite achieving strong performance in both surface reconstruction and novel view synthesis, our method has several limitations that suggest promising directions for future work. First, our model assumes known camera poses as input, which may not always be available in real-world scenarios, especially when the number of input views is extremely sparse. Removing this requirement by exploring pose-free approaches [ye2024no, zhang2025flare] is a compelling future direction that would increase the applicability of our method in unconstrained settings. Moreover, our method assumes a moderate number of input views. For extremely sparse-view settings, the multi-view consistency constraints may become less effective, reducing reconstruction quality. Second, the number of Gaussians grows significantly with the number of input views, potentially limiting the scalability of our method in large-scale or densely captured scenes. Designing more compact or adaptive Gaussian representations, or incorporating efficient pruning and regularization strategies, could help improve scalability without compromising reconstruction quality. Third, our method currently relies on a pre-trained feature extraction network to guide geometric alignment and ensure robustness under varying lighting and viewpoints. This dependence introduces limitations, particularly when the pre-trained features are suboptimal for specific domains or scenes. Exploring more reliable, possibly self-supervised, feature learning strategies or reducing reliance on external networks altogether would further improve robustness and generalization. Addressing these limitations can pave the way for more flexible, efficient, and generalizable Gaussian-based surface reconstruction frameworks.